# Addressing Hallucinations with RAG and NMISS in Italian Healthcare LLM Chatbots

### v2

@@ 4.2 User-Agent Chatbot Framework
This section outlines the architecture of the user-agent chatbot designed to handle queries related to Italian health news, using a RAG approach that integrates a Dense Vector Store (DVS) with LLMs fine-tuned for Italian language processing. This multi-model setup ensures robustness in handling diverse queries related to Italian health news, providing accurate, coherent, and contextually grounded responses.
The system begins with Sentence Transformers to generate dense vector embeddings for the Document Database, which are stored and accessed via the Dense Vector Space. This enables efficient retrieval of relevant information through semantic search and clustering. After experimenting with different chunk lengths, I optimized retrieval performance by selecting a chunk length of 512 tokens with a 64-token overlap, as evaluated by human reviewers on a dataset of 100 health-related queries. For the retrieval part, I choose the Sentence Transformer model ‘paraphrase-multilingual-MiniLM-L12-v2‘ [^1] https://huggingface.co/sentence-transformers/paraphrase-multilingual-MiniLM-L12-v2 for its high-quality embeddings across multiple languages, making it ideal for Italian health news [^2] After testing various chunk token sizes, the chunk length of 512 tokens with a 64-token overlap provide the best balance between retrieval performance and contextual continuity..
Figure 2: User-Agent Application Workflow. Text data is split into chunks for efficient retrieval, indexed using a Sentence Transformer model, and stored as dense vector embeddings in a vector store. The system retrieves relevant chunks based on query embeddings and passes them to an LLM for response generation.
For the retrieval part, I choose various LLMs handling the Italian language. LLaMA2 and LLaMA3 are generative models developed by Meta[49, 30], optimized through supervised fine-tuning (SFT) and reinforcement learning with human feedback (RLHF). Mistral[18] is a 7-billion parameter model offering competitive performance in generating accurate responses. GPT-4[1], the latest multimodal model from OpenAI. Finally, Gemma and Gemma 2[46, 47], lightweight models from Google, are fine-tuned for Italian question-answering and summarization tasks.
The workflow illustrated in Figure 2 starts with retrieving relevant information from a large corpus of chunk news by embedding the user query into a vector space using the sentence transformer. The query embedding is matched against stored dense vector embeddings, and the top-K most similar document chunks are retrieved from the vector database. Next, the generation component uses the retrieved document chunks as context to produce a response. The LLM receives both the original query and the retrieved context, generating a natural language response grounded in factual data, thus reducing hallucinations.
@@ 4.3 Prompt Engineering Template
I develop a question-answering prompting template that guides the model in providing contextually relevant answers based on the information available in the provided context [^3] The list of the final set of Italian questions is available upon request.. An English version is shown in Table 4. This template ensures that the chatbot remains anchored to the information in the provided context. If the answer to a query is not found within the context, the chatbot will respond with: I’m sorry, I can’t help you based on the information I have”.
Table 4: Prompt Engineering - Example (English translation)
Prompt You are a chatbot that loves helping people! Given the following context section, answer the question using only the provided context. If you are unsure and the answer is not explicitly written in the documentation, respond by saying, ’I’m sorry, I can’t help you based on the information I have.’
Question How can myocardial infarction be prevented?
Context [1] The diagnosis of myocardial infarction on-site allows for rapid transportation by equipped ambulance, uncomfortable but not risky, directly to the hemodynamic lab, enabling coronary reopening at least 30 minutes earlier than if the patient arrives at the emergency room with their own means at the reference hospital.” – responsible for myocardial infarction [2] early infarction and saving lives with a simple blood test.” According to a press release from the Mauriziano’s press office. ”In the last ten years,” the press release explains, ”the diagnosis of acute coronary syndromes has progressively shifted attention from the electrocardiogram (always and essential) to the qualitative and quantitative biochemical definition of myocardial damage. The new method provides a more specific, sensitive, and early definition of myocardial damage with undeniable advantages in response speed to the pathology, especially when it presents subtly and unexpectedly. In short, it recognizes infarction practically in real-time.” [3] ”Remembering that acute myocardial infarction (AMI) is a time-dependent pathology,” Mura concludes, ”it is essential to prevent and initiate therapy, either pharmacological or interventional in the hemodynamic room, within the timeframes established by national protocols.
System Based on the context, there is no precise information on the symptoms indicating myocardial infarction. However, it mentions a new hospital technique capable of recognizing infarction in real-time, which is crucial given the time-dependent nature of the pathology.
@@ 5.1 RAG Performance Analysis
The results are presented in Figures 3-4, showing the models’ performances across three levels of questioning. According to [8], a ROUGE score below 0.4 is considered low, between 0.4 and 0.5 is moderate, and above 0.5 is good.
The Gemma model shows the weakest performance across all metrics. Its ROUGE-1 and ROUGE-2 F1 scores fall below 0.4, indicating significant difficulty in capturing relevant unigrams and bigrams. The ROUGE-L score is similarly weak, suggesting issues with structural coherence. Additionally, Gemma’s BLEU and METEOR scores remain low, showing poor n-gram precision and recall. With no EM, Gemma struggles to generate sequences that match the reference.
In contrast, Gemma2 emerges as the top performer across all metrics. Its ROUGE-1 and ROUGE-2 F1 scores exceed 0.5, indicating strong unigram and bigram relevance. The high ROUGE-L score points to solid structural coherence, while the BLEU and METEOR scores suggest excellent n-gram precision and recall. Gemma2 also achieves notable EM rates, ranging between 40% and 50%, highlighting its strength in generating exact matches, which is critical for tasks requiring precise language generation.
GPT-4 delivers balanced performance across all metrics but shows variability across question levels. It follows closely behind Gemma2 in ROUGE, BLEU, and METEOR scores. However, a drop occurs at mid and high question levels, suggesting GPT-4 faces challenges with complex queries. Despite this, its BERTSCORE exceeds 0.9, and it also performs strongly in EM, rivaling Gemma2 in generating exact matches.
Llama2, Llama3, and Mistral exhibit similar patterns. While their F1 scores for all ROUGE variants are relatively low, mainly due to weak precision, their recall scores approach those of Gemma2 and GPT-4. Mistral maintains consistent performance across all question levels, but the Llama models struggle more with more difficult questions.
Overall, Gemma2 emerges as the most performant model, strictly followed by GPT-4, which remains competitive despite a slight decline in performance on complex queries. Mistral and the Llama models perform modestly across ROUGE, BLEU, and METEOR, with Mistral being more stable across all levels. Gemma lags significantly behind, particularly in ROUGE and BLEU. BERTSCORE remains consistently high across all models, with Gemma2 and GPT-4 achieving BERTSCOREs above 0.9.
@@ 5.2 NMISS Performance Evaluation
Results for BLEU and METEOR NMISS scores are shown in Tables 5-6. Findings report the scores against the novel metric for all the models across the question levels. In both cases, NMISS scores generally surpass their standard counterparts, indicating enhanced performance across different difficulty levels.
@@
33footnotetext: Note: Average scores. Higher values are highlighted in bold.
For both metrics, models like GPT-4 show minimal differences between standard and NMISS scores, indicating high consistency in performance across both metrics. However, models like Llama2, Llama3 and Mistral exhibit more significant improvements in their BLEU NMISS scores, suggesting that NMISS better captures the models’ ability to provide additional contextual information, especially in more challenging tasks. Notably, Gemma2 shows strong performance overall, with NMISS scores slightly higher than standard ones, particularly at mid and high levels, highlighting its ability to maintain accuracy and detail in complex scenarios. The increase in NMISS scores when models struggle to answer could be attributed to how NMISS rewards the inclusion of additional relevant information or contextual details. When a model does not provide a concise answer, it often compensates by providing surrounding context, explanations, or related information. While not a perfect match to the expected answer, this additional content contributes meaningfully to the overall response.
@@
33footnotetext: Note: Average scores. Higher values are highlighted in bold.
Tables 7-9 compares ROUGE and ROUGE NMISS scores across different models and difficulty levels. Similar to the trends observed with BLEU and METEOR, the ROUGE NMISS scores generally outperform the standard ROUGE metrics, especially at higher difficulty levels. This consistent pattern across different metrics suggests that NMISS captures additional relevant information, rewarding models for providing more detailed and contextually accurate responses.
Specifically, for most models, ROUGE NMISS scores surpass the standard ROUGE-1 and ROUGE-L scores across all difficulty levels. This is consistent with the behavior observed for BLEU and METEOR. Interestingly, the ROUGE-2 NMISS scores do not consistently outperform the standard ROUGE-2 scores across all models and levels. This deviation from the trend seen in ROUGE-1 and ROUGE-L might be attributed to the focus of ROUGE-2 on bigram overlaps, which are less sensitive to the additional contextual information.
@@
To provide a broad perspective on the various scores, Figures 5-6 illustrate the percentage of instances where NMISS outperforms standard metrics for all questions. For ROUGE-1 and ROUGE-L, NMISS consistently performs better, reflecting its ability to capture more contextually rich and detailed responses. This further highlights NMISS’s effectiveness in evaluating model performance, particularly in greater complexity scenarios.
The performance of ROUGE NMISS for precision across all question levels is as follows: for GPT-4 ROUGE NMISS outperforms ROUGE-1 by approximately 26% at the low level, 30% at the mid level, and 40% at the high level. ROUGE-L shows a similar improvement at all levels, while ROUGE-2 exhibits no notable change. For Llama2 and Llama3, ROUGE NMISS shows remarkable improvement in ROUGE-1, with gains between 80% and 85% across both low and high levels. ROUGE-L demonstrates more modest increases, ranging from 40% to 60%. For Mistral, NMISS provides more modest improvements, ranging from 4% to 65% in ROUGE-1 and similar results for ROUGE-L. Gemma2 follows a similar pattern to GPT-4, with modest improvements across all levels. At the same time, Gemma shows little to no improvement with NMISS, indicating that its issues may stem from machine-related problems rather than contextual relevance.
For recall, similar trends are observed, though with different magnitudes. This is expected, as recall is more sensitive to the length of the context, but they remain noteworthy. GPT-4 gains 18% for ROUGE-1 at mid-level questions, Llama3 improves by 11% at low and 12% at high levels, and Mistral improves by 6% at the low level. Gemma shows no change at the low level but gains 6.1% at mid and high levels.
The same patterns for the models across the question levels apply to METEOR and BLEU percentage analysis, where the difference stands in magnitude. While NMISS heavily impacts BLEU, METEOR is less pronounced.
Figure 5: Percentage of Outperforming of ROUGE NMISS scores
Figure 6: Percentage of Outperforming of BLEU AND METEOR NMISS scores
Overall, the NMISS algorithm demonstrates the greatest improvement at the highest difficulty level, where the gap between NMISS and traditional metrics like BLEU and METEOR is most pronounced. This suggests that NMISS is particularly effective in advanced scenarios that require deeper comprehension and more complex information retrieval or when the model’s predictive power is limited. Notably, with GPT-4 or Gemma2 being the top performers, they benefit the least from NMISS, indicating that answers are already highly contextual and do not need additional elaboration. In contrast, middle-range models like Llama2, Llama3, and Mistral show the most significant gains, as NMISS helps to capture their ability to provide richer contextual information. Gemma sees little to no improvement, indicating that its issues are likely machine-related rather than due to contextual relevance.
These results suggest that when models struggle to provide exact answers, they compensate by offering more detailed contextual information. Importantly, this does not result in hallucinations but rather reflects the models’ ability to maintain accuracy and completeness in the surrounding context. In such cases, even peripheral contexts can contribute to higher NMISS scores, improving the overall response quality. Conversely, where questions are more straightforward, the gap between NMISS and traditional metrics narrows. This indicates that less contextual elaboration is required, and the models perform similarly across both metrics. In summary, NMISS tends to reflect the models’ ability to provide richer, more informative answers in complex situations, which is less evident at lower difficulty levels where responses tend to be more concise and direct, reducing the gap between NMISS and traditional metrics.
6 Conclusions
In this paper, I investigate hallucinations in LLMs using a combination of detection and mitigation techniques. I evaluate a user-agent chatbot interaction with Italian health news articles through a question-answering framework based on the RAG architecture. Models including Gemma, Gemma2 GPT-4, Llama2, Llama3, and Mistral are analyzed. To enhance traditional evaluation metrics like ROUGE, METEOR, and BLEU, I introduce the NMISS scoring system, which emphasizes contextual relevance in model-generated responses. NMISS allows the identification of instances where traditional metrics undervalue contextually rich answers that deviate from the reference.
Regarding performance, Gemma2 and GPT-4 consistently outperformed other models. Gemma2 demonstrated strong results across ROUGE, METEOR, and BLEU, making it a top choice for healthcare-related chatbot applications. GPT-4 excelled in EM accuracy, highlighting its ability to generate sequences closely aligned with reference answers.
The mid-range models Llama2, Llama3, and Mistral show similar performance patterns. Although their F1 scores across all ROUGE variants are lower due to weaker precision, their recall scores are close to those of Gemma2 and GPT-4, indicating their ability to capture a significant amount of relevant content despite generating less precise outputs. Mistral maintains consistent performance across all difficulty levels, while Llama2 and Llama3 struggled with more complex queries.
NMISS application reveals that when a model struggles to provide exact answers, it often compensates by delivering richer contextual information. This does not result in hallucinations but highlights the LLM’s capacity to provide valuable information embedded in the context, even when direct answers are challenging to generate. This demonstrates the importance of context as a carrier of valuable information. NMISS consistently outperforms traditional ROUGE-1 and ROUGE-L scores across all question levels, and similar trends are observed for BLEU and METEOR.
It is important to note that high-performing models like Gemma2 and GPT-4 benefit less from NMISS, which is expected. These models already provide highly precise and contextually accurate responses. In contrast, NMISS proves particularly beneficial for mid-tier models like Llama2, Llama3, and Mistral, which often generate contextually richer responses when faced with more challenging queries. These models benefit from NMISS because it captures and rewards their ability to provide additional relevant context, even if their precision is lower. On the other hand, Gemma shows slight improvement across all metrics, suggesting that its hallucination issues are machine-intrinsic and less related to context-awareness, which NMISS is designed to evaluate.
While RAG models mitigate hallucinations by grounding outputs in external data, NMISS refines the evaluation process. It helps distinguish “false positives” — cases where traditional metrics flag hallucinations due to deviations from the reference, even though the response is contextually accurate. NMISS reduces penalties for models that provide relevant contextual information, even when their answers do not match the reference. This distinction underlines the complementary nature of RAG and NMISS: RAG acts as a tool for preventing hallucinations, while NMISS improves how I evaluate context and relevance in model output, serving as a diagnostic tool for better understanding hallucinations.
Future research could explore extending this framework to more complex datasets, tasks, and languages, further advancing hallucination mitigation and improving LLM performance in real-world applications.
Acknowledgements
This publication was produced while attending the PhD program in Economics and Business at the University of Cagliari, Cycle 38, with the support of a scholarship financed by the Ministerial Decree no. 351 of 9th April 2022, based on the NRRP - funded by the European Union - NextGenerationEU - Mission 4 “Education and Research”, Component 1 “Enhancement of the offer of educational services: from nurseries to universities” - Investment 3.4 “Advanced teaching and university skills”.

### v3

@@ 4.2 Human Evaluation Protocol
To evaluate hallucination detection performance, each model-generated answer was manually annotated as either hallucinated or non-hallucinated as formalized in Section 3.3.2. An answer was labeled as hallucinated if it contained factual inaccuracies, fabricated details, or information not supported by the retrieved context or the corresponding reference answer. This binary annotation served as the ground truth for computing alignment between automatic metrics and human judgment. The annotated dataset forms the basis for the analyses presented in Section 5, enabling direct comparison between traditional metrics and NMISS.
@@ 4.3 User-Agent Chatbot Framework
This section outlines the architecture of the user-agent chatbot designed to handle queries related to Italian health news. The framework adopts a RAG approach that integrates a Dense Vector Store (DVS) with LLMs fine-tuned for Italian language processing. This multi-model setup ensures robustness in managing diverse user queries, providing accurate, coherent, and contextually grounded responses.
Figure 2: User-Agent Application Workflow
Note. Text data is split into chunks for efficient retrieval, indexed using a Sentence Transformer model, and stored as dense vector embeddings in a vector store. The system retrieves relevant chunks based on query embeddings and passes them to an LLM for response generation.
The system begins with a Sentence Transformer model used to generate dense vector embeddings for the document database, which are then stored and indexed within the DVS. This enables efficient retrieval of relevant information through semantic search and clustering. After experimenting with different chunking configurations, retrieval performance was optimized using a chunk length of 512 tokens with a 64-token overlap [^3] Empirically, a chunk size of 512 tokens with a 64-token overlap provided the best trade-off between retrieval accuracy and contextual continuity., evaluated on a dataset of 100 health-related queries.
For embedding generation, I employ the Sentence Transformer model Paraphrase Multilingual MiniLM L12444https://huggingface.co/sentence-transformers/paraphrase-multilingual-MiniLM-L12-v2, selected for its high-quality multilingual embeddings, particularly suitable for Italian health news.
For the generation component, I evaluate multiple LLMs [^5] All open-source models were evaluated with a temperature of 0.7 for comparability. capable of processing Italian text. LLaMA2 and LLaMA3, developed by Meta [touvron2023llama, meta2024llama3], are optimized through Supervised Fine-Tuning (SFT) and Reinforcement Learning with Human Feedback (RLHF). Mistral [jiang2023mistral] is a 7-billion-parameter model known for its efficiency and competitive generation accuracy. GPT-4 [achiam2023gpt] represents the latest multimodal model from OpenAI, while Gemma and Gemma2 [team2024gemma, team2024gemma2] are lightweight models from Google, fine-tuned for question-answering and summarization tasks.
The workflow illustrated in Figure 2 begins by embedding the user query into the same vector space as the preprocessed document chunks. The query embedding is compared against stored vectors to retrieve the top-$K$ most semantically similar chunks from the vector database. These retrieved documents are then supplied to the LLM as contextual input.
During generation, the model receives both the user query and the retrieved evidence, producing a natural language response grounded in factual data. This retrieval-augmented process reduces hallucinations and improves factual coherence across responses.
@@ 4.4 Prompt Engineering Template
I develop a zero-shot question-answering prompting template that guides the model in providing contextually relevant answers based on the information available in the provided context. The structure is shown in Table 4.
Table 4: Prompt Engineering Structure
Prompt Structure
The system prompt instructs the model to behave as a helpful assistant and to generate answers exclusively based on the retrieved context. If the relevant information is not explicitly present in the context, the model is instructed to return a fallback message indicating insufficient information.
Instruction Components
(i) Define the assistant’s role, a chatbot specialized in general healthcare;
(ii) Provide the retrieved context section containing excerpts from the database;
(iii) Ask the model to answer the user’s question using only the provided context;
(iv) Instruct the model to abstain from speculation or external knowledge;
(v) Specify a fallback message for uncertain cases.
This template ensures that the chatbot remains anchored to the information in the provided context. If the answer to a query is not found within the context, the chatbot is instructed to respond with: ”I’m sorry, I can’t help you based on the information I have”.
@@ 5.1 RAG Performance Analysis
The results, summarized in Figures 34, present model performance across three levels of question complexity.
Gemma demonstrates the weakest performance across all evaluation metrics. Its ROUGE-1 and ROUGE-2 F1 scores fall below 0.4666Following dalal2024text, ROUGE scores below 0.4 are considered low, between 0.4 and 0.5 moderate, and above 0.5 good., indicating limited ability to capture relevant unigrams and bigrams. The low ROUGE-L score further suggests poor structural coherence between generated and reference texts. BLEU and METEOR scores remain consistently low, reflecting weak lexical precision and recall, while the absence of EM confirms its difficulty in reproducing reference sequences.
@@
In contrast, Gemma-2 achieves the strongest overall results. Its ROUGE-1 and ROUGE-2 F1 scores exceed 0.5, demonstrating strong lexical alignment, while the high ROUGE-L value indicates superior sentence-level coherence. BLEU and METEOR scores reinforce this trend, highlighting consistent precision and recall advantages. Moreover, Gemma-2 attains the highest exact-match rate, evidencing precise adherence to reference phrasing when contextual alignment permits.
GPT-4 follows closely behind, exhibiting balanced and reliable performance across all metrics. Although slight declines are observed at medium and high difficulty levels, it maintains high overall accuracy, achieving BERTScores above 0.9 and EM values comparable to Gemma-2. These results indicate strong factual grounding and effective adherence to the given prompts.
Llama-2, Llama-3, and Mistral form a consistent mid-tier group. Their ROUGE F1 scores are generally lower, primarily due to reduced precision, though their recall remains comparable to that of the leading models. Among these, Mistral exhibits the most stable performance across all difficulty levels, whereas both Llama models experience sharper performance degradation as question complexity increases.
Overall, a consistent performance hierarchy is observed across all evaluation metrics: Gemma-2 and GPT-4 lead as top performers; Llama-2, Llama-3, and Mistral form a stable middle tier; and Gemma remains the weakest model, often producing responses that lack coherence or contextual grounding. This ranking pattern holds across ROUGE, BLEU, BERTScore and METEOR, indicating robustness in the comparative evaluation.
Among the individual metrics, METEOR yields consistently higher scores than BLEU for all models, as it accounts for stemming, synonymy, and paraphrase matching. This property allows to reward semantically equivalent expressions that BLEU tends to penalize, thereby providing a more lenient yet informative measure of linguistic adequacy. A similar trend is observed for BERTScore, whose values are uniformly high across models due to its semantic sensitivity.
@@ 5.2 NMISS Performance Evaluation
Table 5 reports the percentage of cases in which NMISS outperforms its traditional counterparts, broken down by model and question difficulty. The results highlight NMISS’s added value across evaluation metrics and task complexities.
Across all metrics, Gemma-2 benefits the most from NMISS, achieving the highest outperformance rates despite having one of the lowest hallucination frequencies. This suggests that its outputs, while factually grounded and semantically coherent, often diverge lexically from the reference answers. Conversely, GPT-4 shows smaller NMISS gains, consistent with its strong EM results reported in Section 5.1, indicating close alignment with the reference in both form and structure.
Figures 57 further examine the relationship between hallucination rate and NMISS outperformance across BLEU, METEOR, and ROUGE variants. For BLEU (Figure 5), a clear positive correlation emerges: models with higher hallucination rates, such as LLaMA-2, LLaMA-3, and Mistral, also exhibit greater NMISS improvements. This pattern supports the view that BLEU penalizes elaborative or contextually grounded completions that deviate from reference phrasing, whereas NMISS compensates by rewarding semantic fidelity. Gemma-2 remains an exception, showing substantial NMISS gains despite minimal hallucinations, indicating high-quality, context-enriched responses that standard BLEU scoring undervalues.
Table 5: NMISS Outperformance across Models and Question Levels
Metric Model Low Mid High Overall
BLEU gemma 0.00 0.00 0.00 0.00
BLEU gemma2 87.50 83.33 76.47 82.43
BLEU gpt4 50.00 28.57 40.00 39.52
BLEU llama2 66.67 75.00 63.64 68.43
BLEU llama3 66.67 60.00 73.33 66.67
BLEU mistral 50.00 45.83 52.63 49.49
METEOR gemma 0.00 0.00 0.00 0.00
METEOR gemma2 0.00 21.43 15.00 12.14
METEOR gpt4 0.00 0.00 0.00 0.00
METEOR llama2 0.00 6.25 0.00 2.08
METEOR llama3 4.76 0.00 4.35 3.04
METEOR mistral 0.00 3.70 3.70 2.47
ROUGE-1 F1 gemma 0.00 0.00 0.00 0.00
ROUGE-1 F1 gemma2 81.82 85.71 80.00 82.51
ROUGE-1 F1 gpt4 71.43 28.57 41.18 47.06
ROUGE-1 F1 llama2 75.00 75.00 64.29 71.43
ROUGE-1 F1 llama3 71.43 78.26 73.91 74.53
ROUGE-1 F1 mistral 41.18 57.69 55.56 51.47
ROUGE-2 F1 gemma 0.00 0.00 0.00 0.00
ROUGE-2 F1 gemma2 9.09 7.69 0.00 5.59
ROUGE-2 F1 gpt4 0.00 0.00 0.00 0.00
ROUGE-2 F1 llama2 0.00 0.00 0.00 0.00
ROUGE-2 F1 llama3 0.00 0.00 4.76 1.59
ROUGE-2 F1 mistral 6.67 7.69 0.00 4.79
ROUGE-L F1 gemma 0.00 0.00 0.00 0.00
ROUGE-L F1 gemma2 36.36 50.00 50.00 45.45
ROUGE-L F1 gpt4 0.00 14.29 11.76 8.68
ROUGE-L F1 llama2 66.67 56.25 50.00 57.64
ROUGE-L F1 llama3 52.38 43.48 56.52 50.79
ROUGE-L F1 mistral 29.41 50.00 33.33 37.58
  • Note. Scores represent the percentage of hallucination cases where NMISS outperformed the corresponding traditional metrics, expressed as the ratio of improved cases to valid hallucinations. Bold values highlight the highest scores per metric. “Overall” is the average across all questions difficulty. NMISS outperformance is computed over valid hallucination cases (excluding degenerate scores of 0 or 0.99).
Figure 5: NMISS Utility vs. Hallucination Rate (BLEU)
Figure 6: NMISS Utility vs. Hallucination Rate (METEOR)
Figure 7: NMISS Utility vs. Hallucination Rate (ROUGE Variants)
In the case of METEOR (Figure 6), the effect is less pronounced. This metric already incorporates stemming, synonymy, and paraphrase matching and naturally captures a larger share of semantic variability between system and reference outputs. Consequently, the set of non-matching tokens is comparatively small, reducing the influence of the contextual component in NMISS. Under these conditions, NMISS primarily refines borderline cases rather than producing large-scale adjustments. Nonetheless, Gemma-2 and LLaMA-3 still exhibit meaningful gains, suggesting that their elaborations are contextually grounded even when lexical overlap is already high.
ROUGE analysis (Figure 7) reveals more varied patterns across its variants. ROUGE-1 largely mirrors BLEU’s behavior, with higher hallucination rates correlating with stronger NMISS improvements. ROUGE-L shows a weaker but still positive relationship, while ROUGE-2 remains mostly insensitive to NMISS adjustments, reflecting the rigidity of bigram-based evaluation in capturing nuanced or rephrased content.
Overall, NMISS enhances evaluation in two complementary scenarios: (i) for models that hallucinate more frequently, by distinguishing spurious additions from contextually valid elaborations; and (ii) for high-performing models such as Gemma-2, by recognizing semantically grounded deviations that traditional metrics undervalue. These findings confirm NMISS’s complementary role and robustness across varying task difficulty levels and model qualities. Detailed average NMISS and baseline scores for BLEU, METEOR, and ROUGE are reported in Tables 610 in the Appendix.
6 Conclusions
This study investigates hallucination behavior in LLMs from a dual perspective: mitigation through a RAG architecture and evaluation refinement via the proposed NMISS. The experimental framework focuses on a question-answering task over Italian health news articles, simulating user–agent interactions in a sensitive knowledge domain. Six models—Gemma, Gemma-2, GPT-4, LLaMA-2, LLaMA-3, and Mistral—are evaluated in a zero-shot configuration with a fixed prompt template to ensure consistent generation behavior across systems.
The analysis shows that Gemma-2 and GPT-4 achieve the strongest overall performance under traditional metrics such as BLEU, ROUGE, and METEOR. Gemma-2 demonstrates balanced precision and recall across all difficulty levels, making it a promising candidate for healthcare-related applications. GPT-4, while slightly behind on surface-based metrics, attains the highest EM accuracy, reflecting its strict prompt adherence and literal generation style. Mid-tier models—LLaMA-2, LLaMA-3, and Mistral—achieve comparable recall but lower precision, producing contextually relevant yet lexically diverse completions that traditional metrics tend to undervalue.
To address this evaluative limitation, the NMISS framework is introduced to reward contextual alignment by incorporating information retrieved from the original input document. Focusing on non-hallucinated responses, the analysis quantifies the proportion of cases in which NMISS outperforms standard metrics, yielding several key findings. First, NMISS effectively identifies hidden semantic value in outputs overlooked by conventional metrics. Gemma-2, despite already leading in raw performance, achieves the highest NMISS outperformance across BLEU, ROUGE-1, and METEOR, demonstrating that NMISS can recognize semantically rich yet lexically divergent answers that remain contextually faithful. Second, although GPT-4 performs strongly overall, its NMISS improvement is modest. This reflects its rigid zero-shot generation behavior, which closely mirrors the prompt and minimizes lexical variation, leaving less room for NMISS to detect additional contextual contributions.
Conversely, mid-tier models exhibit substantial NMISS gains, particularly for BLEU and ROUGE-1. Their responses, while more syntactically variable, frequently contain relevant contextual cues that overlap-based metrics penalize but NMISS rewards. This suggests that NMISS captures a compensatory strategy: when these models struggle to respond directly, they often enrich their outputs with contextually grounded elaborations that enhance informativeness.
A positive correlation emerges between hallucination rate and NMISS outperformance for BLEU and ROUGE-1, reinforcing the idea that NMISS excels in borderline cases where literal matching fails to capture valid, context-driven reasoning. Notably, Gemma-2 remains an exception, showing that even factually accurate responses can benefit from a more context-sensitive evaluation.
Overall, the findings demonstrate that while RAG architectures effectively mitigate hallucinations by grounding LLM outputs, NMISS complements this process by refining how output quality is assessed. It compensates for the underestimation introduced by traditional overlap-based metrics, which often assign lower scores to responses that are semantically correct but lexically divergent. By rewarding contextually aligned and informative outputs, NMISS provides a fairer evaluation of model performance. This dual contribution is particularly relevant for high-stakes domains such as healthcare, where contextual accuracy and informativeness often outweigh surface-level similarity.
In conclusion, NMISS represents a valuable addition to the LLM evaluation toolkit. It enhances fairness and interpretability in performance assessment, particularly for models producing rich and context-aware responses. While top-performing models benefit from NMISS refinement, mid-tier models gain even more, as the metric recovers value otherwise overlooked by standard approaches.
Future work could extend NMISS to multilingual and multi-hop question answering, adaptive fine-tuning pipelines, and reinforcement learning frameworks that align model outputs with context-based quality objectives. Beyond the Italian healthcare domain, the combined RAG–NMISS methodology generalizes to other languages and knowledge domains where factual grounding and contextual relevance are essential.
7 Appendix
@@
  • Note: Average scores. Higher values are highlighted in bold.
@@
  • Note: Average scores. Higher values are highlighted in bold.
@@
  • Note: Average scores. Higher values are highlighted in bold.
@@
  • Note: Average scores. Higher values are highlighted in bold.
@@
  • Note: Average scores. Higher values are highlighted in bold.